40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 13 10:50:04 2014, Andres, Update, IMC, Calculation for the input mode cleaner ModeCleanerSetup0.PNGalaModeModeCleanersolution.png
    Reply  Wed Aug 13 13:08:24 2014, rana, Update, IMC, Calculation for the input mode cleaner 
       Reply  Thu Aug 14 15:10:47 2014, Andres, Update, IMC, Calculation for the input mode cleaner SolutionForTheModeCleanerSetup00.pngSolutionForTheModeCleanerSetup11.pngPossibleSetupForModeCleaner.PNGalaModeSolution.zip
    Reply  Wed Aug 13 22:01:57 2014, rana, Update, IMC, Calculation for the input mode cleaner 
Message ID: 10379     Entry time: Wed Aug 13 22:01:57 2014     In reply to: 10374
Author: rana 
Type: Update 
Category: IMC 
Subject: Calculation for the input mode cleaner 

Nic, Andres, and I discussed some more about the MC WFS project today. We want to shorten the proposed WFS2 path. Andres is going to explore moving the 2" diameter lens in coming up with layouts. We also want the WFS to face west so that we can see the diode face with an IR viewer easily and dump the reflected beams in the razor dumps.

We wondered about fixing the power levels and optical gain:

  1. What is the MC modulations depth? What would happen if we increase it a little? Does anyone know how to set it? Will this help the MC frequency noise?
  2. What is the max power on the WFS? I guess it should be set so that the power dissipation of the detector is less than 1 W with the MC unlocked. So P_diss = (100 V)*(I_tot), means that we should have less than 10 mA or ~50 mW when the MC is unlocked.
  3. Another consideration is saturation. The RF signals are tiny, but maybe the DC will saturate if we use any more power. The quadrants are saturated when unlocked and ~200 mV locked. According to D990249, the DC gain in the head is 1000 V/A. The measured power levels going into the heads (w/ MC unlocked) are: P_WFS1 = 4.9 mW and P_WFS2 = 7.7 mW. We don't have control of the DC gain, but there is a 10x and 100x switch available inside the demod board (D980233). From these numbers, I figure that we're in the 100x position and so the effective DC gain between photocurrent and the DC readback voltages is 100 kOhm. Therefore, we are in no danger of optical or electronics saturation. And the unlocked photocurrent of ~40/100000=0.4 mA => 0.04 W heat generated in the diode, so we're OK to increase the power level by another factor of 2-4 if we want.
  4.  We noticed that the ADC inputs are moving by ~50 counts out of 65000, so we're doing a really bad job of signal conditioning. This was previously noticed 6 years ago but we failed to follow up on it. Feh.

While checking this out, I converted the McWFS DC offsets script from csh to bash and committed it to the SVN. We need to remove the prefix 'feature' that Jamie has introduced to cdsutils so that we can use C1 again.

 

ELOG V3.1.3-